The number of non-business bankruptcies has increased over the years. In 2002, the number of non-business bankruptcies was 80,000 less than twice the number in 1994. If the total of non-business bankruptcies for these two years is 2,290,000 find the number of non-business bankruptcies for each year. : The number of non-business bankruptcies has increased over the years. In 2002, the number of non-business bankruptcies was 80,000 less than twice the number in 1994. If the total of non-business bankruptcies for these two years is 2,290,000 find the number of non-business bankruptcies for each year.
Answer by edjones(2391)

let t be the number of bankruptcies in 2002; f number in 1994
t=2f-80,000
t+f=2,290,000
2f-80,000+f=2,290,000
3f=2,370,000
f=790,000
t=1,500,000
.
Ed
Question 148047: The number of non-business bankruptcies has increased over the years. In 2002, the number of non-business bankruptcies was 80,000 less than twice the number in 1994. If the total of non-business bankruptcies for these two years is 2,290,000 find the number of non-business bankruptcies for each year. : The number of non-business bankruptcies has increased over the years. In 2002, the number of non-business bankruptcies was 80,000 less than twice the number in 1994. If the total of non-business bankruptcies for these two years is 2,290,000 find the number of non-business bankruptcies for each year.
Answer by orca(336)

SOLUTION:
Suppose the number of non-business bankruptcies was x in 1994.
Then the number in 2002 was 2x - 80,000
As the total of non-business bankruptcies for these two years is 2,290,000, the equation can be set up as:
x + 2x - 80,000 = 2,290,000
Solving for x, we have
x + 2x = 80,000 + 2,290,000
3x = 2,370,000
x = 790,000
So the number of non-business bankruptcies was 790,000 in 1994, and the number of non-business bankruptcies was 2x - 80,000 = 1,500,000 in 2002.